Please help with an answer. 2/3x-1/6=1/2x+5/6
Answer by thurston(8)

(2/3x)-(1/6)=(1/2x)+(5/6)
Add (1/6) to both sides

(2/3x)=(1/2x)+1
Sub. (1/2) from both sides

(6)(1/6x)=1 (6)
Multiply both sides by six.

x=6
Here is your answer.
Hope you understand it.
Thurston